# Directed Exploration in Reinforcement Learning from Linear Temporal Logic

### v1

@@ 4 Experiments
  • Can DRL2 drive deep exploration in reinforcement learning from linear temporal logic specification?
@@
As the method is designed to handle the full complexity of LTL, our evaluation considers several logic specifications, encompassing reach-avoidance (e.g., $Fa\&G(\neg b)$) and subtask sequences (e.g., $T_{0}:GF(a\&XFc)\&G\neg b$). In order to focus on exploration, the suite of formulas allows easily scaling the number of LDBA states, or the minimum number of steps required in the underlying MDP to induce a transition in the LDBA. To investigate the final question, we perform an evaluation in both tabular and continuous domains. While the former avoids confounding effects arising from function approximation, the latter stresses the ability to scale to complex underlying environments. This also evaluates the versatility of DRL2, as it is in practice coupled with tabular Q-learning [48] and Soft Actor Critic [13], respectively. In the first case, the environment involves navigation in a 2D GridWorld, while in the latter we evaluate dexterous manipulation with a simulated Fetch robotic arm and locomotion of a 12DoF quadruped robot and a 6DoF HalfCheetah. A detailed description of the benchmark environments and specifications is provided in Appendix J; code is available at sites.google.com/view/drl-2.
@@
We compare DRL2 to (1) a counterfactual experience replay scheme (LCER [45]), (2) a novel baseline, that relies on inverse square root visitation counts of LDBA states to compute a potential function for reward shaping (see Appendix J), (3) the underlying learning algorithm with no additional exploration bonuses. We note that other promising approaches for exploration in the LTL domain exist, but they rely on meta-learned components or ad-hoc training regimes [47, 33] and are thus not suitable for a fair comparison.
@@ 4.1 Tabular Setting
Figure 4: Return for Q-learning under eventual discounting. The three rows display reach-avoidance, sequential and circular tasks, as illustrated on the right with optimal policies. Each atomic proposition evaluates to $\top$ in cells of matching color; difficulty increases from left to right. Further details are available in Appendix J. DRL2 is able to drive exploration when naive exploration is insufficient.
We first evaluate the exploration performance of DRL2 by coupling it with tabular Q-learning [48] when operating over discrete state and action spaces in the standard online episodic setting [40]. The evaluation environment is a deterministic 2D gridworld, in which the agent can move in each of the four cardinal directions by one unit at each timestep. The first row of Figure 4 evaluates a reach-avoidance task $G(a\&\neg b)$, in which the agent needs to avoid a large area that covers all but a corridor. The goal area is at the end of said corridor; and the difficulty of the task increases with the length of the corridor. In this case the benefit of DRL2 is evident, as it returns a negative reward whenever the agent leaves the corridor, and the LDBA thus transitions to a sink state. The agent can therefore direct its exploration towards a fraction of the state space, resulting in improved sample efficiency. On the other hand, count-based shaping can only penalize transitions to the sink state once it has been visited enough times. While LCER has the advantage of potentially ignoring failures by hallucinating counterfactual LDBA states during training, it cannot discourage exploration of the forbidden area. We remark that LCER remains a strong method significant when exploration of the underlying MDP is not necessary; a detailed discussion is provided in Appendix E.
@@ 4.2 Continuous Setting
After verifying the effectiveness of DRL2 in interpretable settings, this section investigates if the exploration signal can be scaled to high-dimensional, long-horizon environments requiring the application of deep RL algorithms. For simplicity, we evaluate its application to Soft Actor Critic (SAC) [13], which stands as a fundamental building block for many algorithms [14, 12, 26, 20]. On the top of Figure 5 a simulated Fetch robotic arm [10] is evaluated on two tasks, namely (i) moving its gripper to a specific location while avoiding lateral movements and (ii) gradually producing an horizontal alignment of three cubes. In the middle, an HalfCheetah receives specifications encoding, respectively, finite sequences of positions and infinite sequences of angles for its center of mass, resulting in precise horizontal locomotion, and in front flipping indefinitely. On the bottom, a 12DoF simulated quadruped robot [35] is tasked with (i) fully traversing a narrow corridor, or with (ii) navigating through two zones in sequence. As in the tabular case, the four tasks can all be described through LTL formulas encoding reach-avoidance and sequential behavior. They are reported among further details on the environments in Appendix J.
As expected, in this setting the evaluation is slightly noisier and partially constrained by the learning algorithm. Nevertheless, when applied to significantly more complex underlying environments, DRL2 is competitive with the stronger baselines in simpler problems, and largely outperforms them when exploration in the underlying MDP becomes more challenging. Interestingly, we observe that counterfactual experience replay (LCER) is less effective in this setting. We remark that LCER can generate unfeasible states for the product MDP, which can be harmful when the agent has the ability to interpolate between training samples.
@@
Figure 5: Return for SAC under eventual discounting on Fetch (left), Doggo (center) and HalfCheetah (right), as shown in renderings. DRL2 confirms its ability to drive exploration in complex tasks when coupled with deep RL.
@@ 5 Related Work
Learning from LTL specification has seen remarkable progress in recent years. While this section provides an essential overview, an extended selection of works is presented in Appendix H.
LTL is among several languages for task specification in RL: for instance, numerous works have investigated Reward Machines [21, 22], which however do not match the expressiveness of LTL. While Reward Machines can be very effective for reward shaping, either through heuristics [6] or value iteration [7], these methods are static and do not naively generalize to our setting.
@@
The combination of LTL and RL has initially focused on reconciling logic and policy optimization through the definition of a product MDP and the design of a reward signal encouraging task satisfaction [6, 28, 17, 7, 18, 4, 24]. This led to the development of principled approaches, proposing schemes that provably and directly optimize a lower bound on the likelihood of formula satisfaction [45, 37].
@@
The sparsity of rewarding schemes has motivated the development of several methods, involving the use of heuristics [28], an accepting frontier function [17, 18], bonuses for first visitations [5], or the knowledge of additional information such as annotated maps [49]. Recently, Shah et al. [36] also propose a heuristic scheme for shaping constraints expressed in LTL. This scheme is partially aligned with the shaping signal that DRL2 naturally recovers, but only considers accepting paths and is restricted to on-policy methods. Directly within the LTL literature, another line of work leverages the discrete structure of automata and learns hierarchical [18, 23, 22], goal-conditioned [33] or modular [5] policies. While considerably improving sample efficiently, these approaches are known to potentially introduce suboptimality [22].
@@
A further possibility is that of leveraging control over the automaton to produce synthetic experience by relabeling the LDBA states of collected transitions. Such schemes have been proposed in the context of Reward Machines [22] and LTL [45]; as they can produce off-policy samples, they cannot be naively applied to on-policy methods, as discussed in Appendix E. Other relabeling techniques only target the initial LDBA state [46] or focus on optimality guarantees [37].
Finally, we note that previous work have also led to strong exploration on unseen tasks, as the result of a meta-learning phase when a distribution over MDPs can be accessed [43, 27].
@@ 6 Discussion
DRL2 is not designed to directly address certain exploration issues (such as $\epsilon-$transitions). Nonetheless, it can be seamlessly combined with existing experience replay methods that do. As several other exploration schemes, we note that DRL2 introduces a slight non-stationarity in the reward signal, which needs to be addressed by the underlying learning algorithm.
@@
This work opens up exciting future directions, including a formal analysis of which reward shaping terms would not only grant consistency, but also maximize sample efficiency. Moreover, as DRL2 leverages the LDBA structure, its effectiveness is dependent on it. Its applicability can thus further benefit by LDBA construction algorithms that do not return an arbitrary LDBA in its equivalence class, but rather the one which is most suitable for guiding an RL agent.
Having shown how the fundamental problem of sparsity for complex LTL tasks can be addressed by directly leveraging their structure, we believe that the evidence provided in this work further supports reinforcement learning as suitable paradigm for extracting a controller from a logic specification, simply via interaction and learning.
@@ Acknowledgments
As stated above, this invariance relies on the assumption that the product of discounts converges to zero, which would happen in case an accepting state is visited infinitely often. Intuitively, once a good policy is found, the contribution of the intrinsic reward fades, and there is no incentive for optimization algorithms to update the policy away from its optimum. While this assumption often holds for the optimal policy (e.g., in deterministic environments), there are cases in which this may not be true (e.g., in the stochastic case, if the LDBA features a sink state). Nevertheless, we note that the assumption holds for optimal policies in tasks considered in this paper, and in recent literature [45].
@@ Appendix C Addition of a Virtual Sink State
In order to provide an informative learning signal, DRL2 augments LDBAs that do not feature a sink state with a virtual one and assumes that it can be reached from any other LDBA state. We note that this is equivalent with associating each LDBA state with a discount factor equal to the likelihood of transitioning to the virtual sink. While this can introduce myopic behavior [45], we remark that DRL2 also learns and updates its estimate of the LDBA transition kernel. As any virtual sink state cannot in practice be reached, the likelihood of transitioning to it from each visited LDBA state converges to zero as the number of environment steps increases. Thus, as training progresses, any virtual sink is effectively removed, together with any degree of myopia introduced.

### v2

@@ 4 Experiments
  • Can DRL2 drive deep exploration in reinforcement learning from LTL specification?
@@
As the method is designed to handle the full complexity of LTL, our evaluation considers several logic specifications, encompassing reach-avoidance (e.g., $Fa\&G(\neg b)$) and subtask sequences (e.g., $T_{0}:GF(a\&XFc)\&G\neg b$). In order to focus on exploration, the suite of formulas allows easily scaling the number of LDBA states, or the minimum number of steps required in the underlying MDP to induce a transition in the LDBA. To investigate the final question, we perform an evaluation in both tabular and continuous domains. While the former avoids confounding effects arising from function approximation, the latter stresses the ability to scale to complex underlying environments. This also evaluates the versatility of DRL2, as it is in practice coupled with tabular Q-learning (Watkins & Dayan, 1992) and Soft Actor Critic (Haarnoja et al., 2018a), respectively. In the first case, the environment involves navigation in a 2D gridworld, while in the latter we evaluate dexterous manipulation with a simulated Fetch robotic arm and locomotion of a 12DoF quadruped robot and a 6DoF HalfCheetah. A detailed description of the benchmark environments and specifications is provided in Appendix M; code is available at github.com/marbaga/drl2.
@@
We compare DRL2 to:
  1. 1.
    a counterfactual experience replay scheme (LCER (Voloshin et al., 2023)),
  2. 2.
    a novel count-based baseline  [^4] This baseline is, to the best of our knowledge, novel, as the application of standard RL exploration ideas to the MRP/MDP derived from the LDBA has not been thoroughly investigated. We choose a count-based approach as the MRP is discrete, and does not require hashing or continuous generalizations Burda et al. (2019), while a straightforward application of reward-based optimistic methods is not possible, as the reward of the MRP is known. , that relies on inverse square root visitation counts of LDBA states to compute a potential function for reward shaping (Tang et al., 2017). If $n(b)$ is the visitation count of LDBA state $b\in\mathcal{B}$ from the beginning of training, the intrinsic reward for the high-level transition $b\to b^{\prime}$ is simply computed as $R_{\text{intr}}(b,b^{\prime})=\gamma^{\times}(b^{\prime})\frac{1}{\sqrt{n(b^{\prime})}}-\frac{1}{\sqrt{n(b)}}$.
  3. 3.
    the underlying learning algorithm with no additional exploration bonuses; this corresponds to simply setting $R_{\text{intr}}(b,b^{\prime})=0$ in DRL2.
All shared hyperparameters are set to the ones maximizing the performance of the No exploration baseline; the prior strength $\alpha$ is the main hyperparameter introduced by DRL2, and is simply set to $10^{3}$ in all tabular experiments, and $10^{5}$ in all continuous experiments. For a thorough discussion on hyperparameters, we refer to Appendix M. We note that other promising approaches for exploration in the LTL domain exist, but they rely on meta-learned components or ad-hoc training regimes (Wang et al., 2023, Qiu et al., 2023) and are thus not suitable for a fair comparison.
@@ 4.1 Tabular Setting
Figure 4: Return for Q-learning under eventual discounting. The three rows display reach-avoidance, sequential and circular tasks, as illustrated on the right with optimal policies. Each atomic proposition evaluates to $\top$ in cells of matching color; difficulty increases from left to right. Further details are available in Appendix M. DRL2 is able to drive exploration when naive exploration is insufficient.
We first evaluate the exploration performance of DRL2 by coupling it with tabular Q-learning (Watkins & Dayan, 1992) when operating over discrete state and action spaces in the standard online episodic setting (Sutton & Barto, 2018). The evaluation environment is a deterministic 2D gridworld, in which the agent can move in each of the four cardinal directions by one unit at each timestep. The first row of Figure 4 evaluates a reach-avoidance task $G(a\&\neg b)$, in which the agent needs to avoid a large area that covers all but a corridor. The goal area is at the end of said corridor; and the difficulty of the task increases with the length of the corridor. In this case the benefit of DRL2 is evident, as it returns a negative reward whenever the agent leaves the corridor, and the LDBA thus transitions to a sink state. The agent can therefore direct its exploration towards a fraction of the state space, resulting in improved sample efficiency. On the other hand, count-based shaping can only penalize transitions to the sink state once it has been visited enough times. While LCER has the advantage of potentially ignoring failures by hallucinating counterfactual LDBA states during training, it cannot discourage exploration of the forbidden area. We remark that LCER remains a strong method significant when exploration of the underlying MDP is not necessary; a detailed discussion is provided in Appendix E.
@@ 4.2 Continuous Setting
After verifying the effectiveness of DRL2 in interpretable settings, this section investigates if the exploration signal can be scaled to high-dimensional, long-horizon environments requiring the application of deep reinforcement learning algorithms. For simplicity, we evaluate its application to Soft Actor Critic (SAC) (Haarnoja et al., 2018a), which stands as a fundamental building block for many algorithms (Haarnoja et al., 2018b, Eysenbach et al., 2019, Kumar et al., 2020, Ibarz et al., 2021). On the top of Figure 5 a simulated Fetch robotic arm (de Lazcano et al., 2023) is evaluated on two tasks, namely (i) moving its gripper to a specific location while avoiding lateral movements and (ii) gradually producing an horizontal alignment of three cubes. In the middle, an HalfCheetah receives specifications encoding, respectively, finite sequences of positions and infinite sequences of angles for its center of mass, resulting in precise horizontal locomotion, and in front flipping indefinitely. On the bottom, a 12DoF simulated quadruped robot (Ray et al., 2019) is tasked with (i) fully traversing a narrow corridor, or with (ii) navigating through two zones in sequence. As in the tabular case, the four tasks can all be described through LTL formulas encoding reach-avoidance and sequential behavior. They are reported among further details on the environments in Appendix M.
@@
Figure 5: Return for SAC under eventual discounting on Fetch (left), Doggo (center) and HalfCheetah (right), as shown in renderings. DRL2 confirms its ability to drive exploration in complex tasks when coupled with deep reinforcement learning.
As expected, in this setting the evaluation is significantly noisier and partially constrained by the learning algorithm. Nevertheless, we observe that DRL2 is competitive with the stronger baselines in simpler problems, and likely to outperform them when exploration in the underlying MDP becomes more challenging. Interestingly, we observe that counterfactual experience replay (LCER) is less effective in this setting. We remark that LCER can generate unfeasible states for the product MDP, which can be harmful when the agent has the ability to interpolate between training samples.
@@ 5 Related Work
Logic for Task Specification in RL
LTL is among several logic languages for task specification in reinforcement learning: for instance, numerous related works have investigated Reward Machines (Icarte et al., 2018, Camacho et al., 2019, Vaezipoor et al., 2021, Icarte et al., 2022). While Reward Machines can handle regular expressions, LTL is designed to address a superset, namely $\omega$-expressions. As a consequence, LTL gains the ability to reason about infinite sequences and tasks including liveness and stability.
@@
For this reason, several works, including our own, have investigated LTL for task specification. Early attempts focused on reconciling logic and policy optimization through the definition of a product MDP and the design of a reward signal encouraging task satisfaction (Camacho et al., 2017, Li et al., 2017, Hasanbeig et al., 2018, Camacho et al., 2019, Hasanbeig et al., 2020, Bozkurt et al., 2020, Kantaros, 2022). Several works relied on heuristic reward shaping schemes (Li et al., 2017), which gradually evolved towards dynamic or eventual rewarding and discounting schemes (Hasanbeig et al., 2018; 2020, De Giacomo et al., 2020b; a, Cai et al., 2021, Voloshin et al., 2023). This led to the development of principled approaches, proposing schemes that provably and directly optimize a lower bound on the likelihood of formula satisfaction (Voloshin et al., 2023, Shao & Kwiatkowska, 2023).
As for the optimization of such schemes, while linear programming is sufficient for known dynamics and finite action spaces, Q-learning approaches have been adopted when a model of the environment is not available (Hasanbeig et al., 2018, Bozkurt et al., 2020, Cai et al., 2021). A fundamental restriction of this approaches was lifted by adopting actor-critic (Hasanbeig et al., 2020) or policy gradient (Voloshin et al., 2023) methods suited for continuous action spaces. Nevertheless, in this case, empirical demonstration of LTL-guided reinforcement learning have remained mostly constrained to low-dimensional setting.
This work builds on top of these efforts, and adopts established rewarding and discounting schemes. We focus, however, on the specific problem of exploration, and propose a method to densify the reward signal. Furthermore, leveraging modern actor-critic methods, we extend our empirical evaluation to high-dimensional, complex environments.
Exploration in Logic-driven Reinforcement Learning
@@
A significant effort has been targeted at addressing the sparsity arising from existing rewarding schemes (Hasanbeig et al., 2018, Voloshin et al., 2023), involving the use of heuristics (Li et al., 2017), an accepting frontier function (Hasanbeig et al., 2018; 2020), bonuses for first visitations (Cai et al., 2021), or the knowledge of additional information such as annotated maps (Xiong et al., 2023). Recently, Shah et al. (2024) also propose a heuristic scheme for shaping constraints expressed in LTL. This scheme is partially aligned with the shaping signal that DRL2 naturally recovers, but only considers accepting paths and is restricted to on-policy methods. Directly within the LTL literature, another line of work leverages the discrete structure of automata and learns hierarchical (Hasanbeig et al., 2020, Jothimurugan et al., 2021, Icarte et al., 2022), goal-conditioned (Qiu et al., 2023) or modular (Cai et al., 2021) policies. While considerably improving sample efficiently, the approaches listed so far are known to potentially introduce suboptimality (Icarte et al., 2022). In contrast, DRL2 recovers the optimal policy under mild assumptions, as described in Appendix A.
In the parallel direction of Reward Machines, several works investigate reward shaping for reinforcement learning from logic specification. Camacho et al. (2017) consider potential functions $\Phi(s,b)$ of both the MDP and automaton states; among those, they investigate counting the minimum number of steps to reach an accepting automaton state, which is equivalent to computing values under a specific automaton transition kernel. Another work (Icarte et al., 2022) explores an extension, which is also fundamentally related to our method: the authors investigate value iteration on an MDP synthetised from the automaton to compute a potential function for reward shaping, thus leveraging the semantics of the task. While motivated by a similar intuition to ours, a naive application of this method would not be viable in an eventually discounted setting, which involves $\omega$-regular problems, as we report in Appendix F. By adopting an LDBA instead of a DFA, and more importantly by integrating an eventual discounting scheme, our method is able to avoid myopic behavior while handling the full complexity of LTL. Futhermore, DRL2 does not require assumptions on controllability by defining a Markov reward process instead. Values estimated in the MRP are thus softer than the optimal ones in a corresponding MDP, and we find them to be more informative for exploration. We validate this discussion empirically, by evaluating optimal transition kernels under eventual discounting in Appendix F. By updating its distribution over MRP transition kernels, it is moreover capable of providing an adaptive and informative reward signal, rather than a static one derived from a hard assumption.
@@
Beyond reward shaping, some works leverage control over the automaton to produce synthetic experience by relabeling the LDBA states of collected transitions. Such schemes have been proposed in the context of Reward Machines (Icarte et al., 2022) and LTL (Voloshin et al., 2023); as they can produce off-policy samples, they cannot be naively applied to on-policy methods, as discussed in Appendix E. Other relabeling techniques only target the initial LDBA state (Wang et al., 2020) or focus on optimality guarantees (Shao & Kwiatkowska, 2023). These methods are essentially orthogonal to DRL2 (see Appendix E), and may be combined. Finally, we note that previous work have also led to strong exploration on unseen tasks, as the result of a meta-learning phase when a distribution over MDPs can be accessed (Vaezipoor et al., 2021, León et al., 2022).
@@
Directed Exploration in Reinforcement Learning
In this work, we use the word directed with a specific meaning, as DRL2 directs the agent towards accepting LDBA states, by rewarding LDBA transitions towards them. In contrast, count-based methods promote transitions to any unvisited LDBA state, and may thus be considered undirected. Nonetheless, this idea has parallels in non-logic-conditioned reinforcement learning, as a direction can be given with respect to entities beyond accepting states. In this context, directed has been used in conjunction to the Bayesian notion of information, in order to promote exploration of states that are maximally informative of some quantity (e.g., a reward (Lindner et al., 2021), a value estimate (Nikolov et al., 2019), or learned dynamics (Guo & Brunskill, 2019)).
Another way to provide direction or guidance is in the context of action-masking methods (Stolz et al., 2024, Theile et al., 2024), which pose hard constraints on actions that can be selected by the agent. DRL2 provides a milder feedback through an intrinsic reward term, and does not explicitly prevent actions from being taken. Finally, many works have explored goal-conditioned RL as a technique to guide and improve exploration (Eysenbach et al., 2022), with several works at the intersection with LTL (Qiu et al., 2023, Yalcinkaya et al., 2024). Similarly to goal-conditioned RL, our work extracts a reward from a specific task (in our case, a logic-formula); however, this task is unique, and we do not consider the standard multi-goal/universal value function formulation (Schaul et al., 2015) that is at the core of goal-conditioned RL.
@@ 6 Discussion
DRL2 is generally capable of improving exploration in LTL-conditioned RL; however, we do not expect it to be useful for all tasks, and we suggest four such instances. First, while DRL2 reduces sparsity in reward signal distilled from LTL specifications, the shaped reward is not entirely dense. Intuitively, instead of receiving a non-zero reward when an accepting state is visited, the agent receives it at every LDBA transition. However, if naive exploration is insufficient to induce an LDBA transition, exploration remains challenging. For this reason, we expect our method to be insufficient in environments with extremely long horizons and rare LDBA transitions. Second, in tasks in which each MDP state can be associated with many LDBA states, or presenting $\epsilon$-transitions, a replay method such as LCER is often sufficient, and, while it can be seamlessly combined, DRL2 may not provide significant improvements (see Appendix E). Third, DRL2 relies on an symmetric prior over LDBA transitions, which is in general effective, but could be misleading for specific tasks and MDP morphologies. Fourth, as several other exploration schemes, we note that DRL2 introduces a slight non-stationarity in the reward signal, which needs to be addressed by the underlying learning algorithm and might be problematic in very stochastic environments, especially for off-policy algorithms. Nevertheless, this is a common issue of intrinsic exploration methods, which is in practice often neglectable Burda et al. (2019). Finally, and perhaps foremost, DRL2 relies on the availability of a correct LDBA, and a ground-truth, accurate labeling function to infer AP evaluations from states. While this assumption is shared by the bulk of works in LTL-conditioned RL, real-world deployment of such algorithms would require algorithms that are robust to partial observability and noise, and capable of jointly inferring the labeling function from data, which we highlight as an important research direction.
@@
This work opens up exciting future directions, including a formal analysis of which reward shaping terms would not only grant consistency, but also maximize sample efficiency. Moreover, as DRL2 leverages the LDBA structure, its effectiveness is dependent on it. Its applicability can thus further benefit by LDBA construction algorithms that do not return an arbitrary LDBA in its equivalence class, but rather the one which is most suitable for guiding a reinforcement learning agent.
Conclusion
This work addresses the fundamental problem of sparsity for complex LTL tasks by directly leveraging their structure. Thus, we believe that the evidence provided in this work further supports the effectiveness of reinforcement learning as a paradigm for extracting a controller from a logic specification.
@@ Acknowledgments
As stated above, this invariance relies on the assumption that the product of discounts converges to zero, which would happen in case an accepting state is visited infinitely often  [^5] Formally, convergence to zero only happens if the agent visits an accepting state infinitely many times over an infinite horizon, which occurs if and only if the policy satisfies the formula. In practice, $\Gamma_{t}$ decays exponentially fast as the number of visits until $t$ increases: if $\gamma=0.99$ and an accepting state has been visited, say, $1000$ times, then $\Gamma_{t}\approx 10^{-4}$. . Intuitively, once a good policy is found, the contribution of the intrinsic reward fades, and there is no incentive for optimization algorithms to update the policy away from its optimum. While this assumption often holds for the optimal policy (e.g., in deterministic environments), there are cases in which this may not be true (e.g., in the stochastic case, if the LDBA features a sink state). Nevertheless, we note that the assumption holds for optimal policies in tasks considered in this paper, and in recent literature (Voloshin et al., 2023).
@@ Appendix C Addition of a Virtual Sink State
In order to provide an informative learning signal, DRL2 augments LDBAs that do not feature a sink state with a virtual one and assumes that it can be reached from any other LDBA state. We note that this is equivalent with associating each LDBA state with a discount factor equal to the likelihood of transitioning to the virtual sink. While this can introduce myopic behavior (Voloshin et al., 2023), we remark that DRL2 also learns and updates its estimate of the LDBA transition kernel. As any virtual sink state cannot in practice be reached, the likelihood of transitioning to it from each visited LDBA state converges to zero as the number of environment steps increases. Thus, as training progresses, any virtual sink is effectively removed, together with any degree of myopia introduced.